Find the soultion to the system of equations.
4x + 1 over 2y = -12
3x + 2y = 4

Answer by checkley77(12844)

4x + y/2 = -12 multiply this equation by -4 & add
3x + 2y = 4
-16x-2y=48
------------------
-13x=52
x=52/-13
x=-4 ans.
4*-4+y/2=-12
-16+y/2=-12
y/2=-12+16
y/2=4
y=2*4
y=8 ans.
Proof:
3*-4+2*8=4
-12+16=4
4=4